OpenStudy (anonymous):

solve: (step by step if you would) s^3 + 4s^2 - s - 4 = 0

OpenStudy (anonymous):

Look at Part 2 at http://www.wikihow.com/Factor-a-Cubic-Polynomial We know that s³ + 4s² - s - 4 = (s+1)(s-1)(s+4) = 0 So s = 1, -1, and -4

OpenStudy (precal):